OpenStudy (anonymous):

The point (3, -4) is on the terminal side of an angle theta . What is cos theta ?

OpenStudy (anonymous):

@tcarroll010 u got this

OpenStudy (anonymous):

 remember=

OpenStudy (anonymous):

cosine is the x-coordinate divided by the length from (0,0) to (3, -4)

OpenStudy (anonymous):

cos (-x) = cos (x) So, you can go with @dpalnc 's drawing or make that -4 a +4 and just look at the first quadrant. That's what I do, but different people look at things differently. And since the cos is: adjacent over hypotenuse you look for the "x" value over the lenght of the hypotenuse and you use the Pythoagorean Theorem to get the length: a^2 + b^2 = c^2 You already know "a" and "b".